(3)A 100% concetrate is to be mixed with a mixture having a concentrate of 40% to obtain 55 gallons of a mixture with a concentrate of 75%. how much of the 100% concentrate will be needed?

Answer by checkley75(3666)

1X+.4(55-X)=.75*55
X+22-.4X=41.25
.6X=41.25-22
.6X=19.25
X=19.25/.6
X=32.0833 GALLONS OF 100% SOLUTION IS NEEDED.
PROOF
32.0833+.4(55-32.08333)=41.25
32.08333+22-12.8333=41.25
54.08333-12.8333=41.25
41.25=41.25